OpenStudy (anonymous):

How many real solutions does the equation x^2 + 4x - n^2 = 0 have? A.None B. One C. Two D. Cannot be determined Also, we are learning how to choose "Cannot be determined" properly. So make sure 100% that you can't solve it, before answering cannot be determined! Thanks

OpenStudy (anonymous):

C. Two

OpenStudy (anonymous):

Explanation?

Parth (parthkohli):

This is a quadratic equation. Apparently, all the equations with the degree as '2' have two solutions.

OpenStudy (anonymous):

@CoCoTsoi will love to give tat!

OpenStudy (shubhamsrg):

discriminant = b^2 - 4ac..here its 16+4n^2 now here assuming n is a real no,,can you tell me whether discriminant is greater than 0,less than it or maybe equal to it.?

OpenStudy (apoorvk):

To check if a quadratic has real factors (solutions) or not, check the Discriminant "D" \[D = b^2 - 4ac\] where a,b and ccan be found by comparing your quadratic with the standard form: \[ax^2 + bx +c = 0\] now, if: D=0. --->One real root or factor (that means its a complete square) D>0 --->Two real factors D<0 ---> NO real factors So, what do you think now for the above problem?

OpenStudy (anonymous):

For some reason I got -15... which according to @apoorvk is no real factors, something must be wrong hmmm

OpenStudy (apoorvk):

Answer is C. Why, let me show you First tell me this - what would a,b,c in this case be?

OpenStudy (anonymous):

i worked it again. For some reason I thought the original equation wasn't in standard form (is it?).. I got 20 this time.

OpenStudy (anonymous):

20 (D) > 0, so would it be c?

OpenStudy (anonymous):

a would be 1, b would be 4, and c would be -1? @apoorvk

OpenStudy (shubhamsrg):

how can you get 20 or some other constant when you dont know n..

OpenStudy (apoorvk):

c would be "-n^2"

OpenStudy (apoorvk):

think..

OpenStudy (shubhamsrg):

c= -n^2 .. just treat it like any other constant real no..

OpenStudy (anonymous):

@shubhamsrg well I assumed if there is a - sign there is a 1 in front of n? is that stupid?

OpenStudy (anonymous):

Are there any conditions for n?

OpenStudy (apoorvk):

no that isn't stupid. but the whole constant term is "-n^2" right? that 'n^2' is definitely not a variable

OpenStudy (shubhamsrg):

well n is not a variable and as i said before..treat is like any other constant no.

OpenStudy (anonymous):

ok...

OpenStudy (apoorvk):

I guess too many cooks are spoiling the broth here, so should I leave it to you guys at the moment?

OpenStudy (anonymous):

No no no.. I still need help from everyone. I'm still a little confused

OpenStudy (apoorvk):

okay, np

OpenStudy (anonymous):

x^2+4x-n^2=0 4^2-4(1)(-n^2) =16+4n^2 if there is two roots, 16+4n^2>0 4n^2>-16 n^2 >-4 n>2i or n<-2i if there is one root, 16+4n^2=0 n=2i or -2i if there is no root, 16+4n^2<0 -2i<n<2i that means if n can be all values, we can get different numbers of roots. I would choose cannot determine

OpenStudy (shubhamsrg):

@CoCoTsoi what is say is true..but at this standard i dont think they want us to consider n as a complex no..

OpenStudy (shubhamsrg):

i mean what you say is true *

OpenStudy (anonymous):

That's why I asked if there is condition for n :P

OpenStudy (anonymous):

keep in mind this is an algebra 1 class

OpenStudy (anonymous):

I would still choose cannot be determined as we do not know what n is :D

OpenStudy (anonymous):

@apoorvk You think it's c? Can you show me the steps? I'm kinda lost? Who else things it's c, two solutions?

OpenStudy (apoorvk):

Just tell me something, Kiera - just soppose that 'n' is any constant. so n62 would be any constant - right? In place of 'c' in ax^2 +bx +c in your equation, we have '-n^2' right?

OpenStudy (apoorvk):

*suppose and & **n^2

OpenStudy (anonymous):

yes

OpenStudy (apoorvk):

So what would the discriminant be?

OpenStudy (anonymous):

ok what would the c be in the discriminant?

OpenStudy (anonymous):

would it be like -1 or just -n?

OpenStudy (anonymous):

@Rohangrr copied the answer from here: http://answers.yahoo.com/question/index?qid=20100630112659AA1OiUz

OpenStudy (apoorvk):

D = b^2 - 4ac = 4^2 - 4(1)(-(n^2)) = 16 + 4n^2 ---> am i right?

OpenStudy (callisto):

@apoorvk Yes

OpenStudy (apoorvk):

@FoolForMath nothing new with that - Y! Answers is @Rohangrr 's favorite. @Callisto i was asking the kid - lol.

OpenStudy (callisto):

@KieraJan Have you learnt complex number yet?

OpenStudy (apoorvk):

@ @KieraJan suppose the equation would have been "x^2 + 4x - 6" - what would have the 'c' be been?

OpenStudy (apoorvk):

no @Callisto we are making it too complicated - this is really a stupid problem - there is no complex no. involved.

OpenStudy (anonymous):

@Callisto no

OpenStudy (apoorvk):

@KieraJan - I asked - suppose the equation would have been "x^2 + 4x - 6" - what would have the 'c' been in this case??

OpenStudy (anonymous):

D = b^2 - 4ac = 4^2 - 4(1)(-(n^2)) = 16 + 4n^2 - I followed that @apoorvk

OpenStudy (anonymous):

-6 right?

OpenStudy (apoorvk):

very well - that means we are done!! well almost.. so now tell me - can n^2 be negative ever?

OpenStudy (anonymous):

I'm not sure...

OpenStudy (anonymous):

i guess it was in this problem?

OpenStudy (apoorvk):

No, don't guess - how did you arrive at '-6'?

OpenStudy (anonymous):

because there was a negative in front of 6, the c.. so -6

OpenStudy (anonymous):

so if there is a negative in front of n^2, I'm thinking it could probably be a negative

OpenStudy (apoorvk):

Gorget '-6'. A whole square can NEVER be negative. because a positive no. multiplied by itself is always positive, as well as a negative no. multiplied by itself is postive - because two '-' signs make a plus sign - right? for g.--> (-4)^2 = -4 x -4 = +16

OpenStudy (anonymous):

yes you are right,

OpenStudy (anonymous):

yes because to negatives make a positive

OpenStudy (apoorvk):

yeah. but the whole equation was---> 4^2 - (- 4xn^2) --> right?? adjacebt minus sign makes a plus right? so it transforms to --> 16 '+' 4n^2 - right?

OpenStudy (anonymous):

yes

OpenStudy (apoorvk):

that means that, whatever 'n' is - negative OR positive, n^2 will always be positive - right?

OpenStudy (anonymous):

right!

OpenStudy (apoorvk):

and if n^2 is positive, then 4n^2 would also be positive - positive multiplied by a positive is positive I heard. - okay?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so we have 16 + 4n^2

OpenStudy (apoorvk):

and now, 16 added to 4n^2, which is "16 + 4n^2" is also positive - since positive added to positive is positive as well?

OpenStudy (apoorvk):

yes, ,we have 16+4n^2, which is the DISCRIMINANT, which is positive - ALWAYS!

OpenStudy (anonymous):

and you said D>0 is two solutions.. so how does that tie in to it?

OpenStudy (apoorvk):

the minimum - D can be in this case is when n=0, that is D is atleast 16. so always greater 0 anyhow - so two solutions.

OpenStudy (anonymous):

so to show my work could I do: 16 + 4n^2 = 20n^2 which 20n^2 > 0 , therefor two solutions (answer choice C) @ apoorvk

OpenStudy (apoorvk):

NO!! 16 + 4n62 is that itself - how can you add two terms '16" and "4n^2"? does "2x + 4' equal "6x"? NO - similarly here as well.

OpenStudy (anonymous):

so I could say 16 + 4n^2 > 0 , therefor two solutions?

OpenStudy (apoorvk):

**(16 + 4n^2) You just let it remain that way. As I said each term is always positive - "16" is positive. "4n^2" is positive - always - as we derived it above^. so positive plus positive is also positive. hence the whole "16 + 4n62" is positive too.

OpenStudy (apoorvk):

yeah.

OpenStudy (anonymous):

so either way since neither is negative, even when added together, it won't go below 0, therefor it will always be two solutions, greater than 0?

OpenStudy (apoorvk):

Exactly!!!! Now you're getting it!

OpenStudy (anonymous):

ah ha! (:

OpenStudy (anonymous):

so if you were doing my homework, you would agree that it was C, two solutions? and not Cannot be determined?

OpenStudy (apoorvk):

hundred percent.

OpenStudy (anonymous):

good!

OpenStudy (anonymous):

So I now solved the problem completely, I think?

OpenStudy (apoorvk):

yup - just follow up with the correct statements - you have done that I hope?

OpenStudy (anonymous):

Thanks to everyone for your help! (: yes I have

OpenStudy (apoorvk):

very good. no worries :)